OpenStudy (anonymous):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. My equation is f(x)=4x↑2-48x+156 Thanks in advance

OpenStudy (anonymous):

f(x)=4x\(^2\)-48x+156 substituting x for 4 gives f(4) = 4(4)\(^2\)-48(4)+156 = 28 and for 7 gives f(7) = 4(7)\(^2\)-48(7)+156 = 23 so between 7 and 4 there's a 23-28 = -5 difference 7 and 4 are 7-4=3 spaces apart, so the average rate of change is -5/3

OpenStudy (anonymous):

Thank you so much- I understand it much better now. Thanks!